# Demystifying Workload Imbalances in Large Transformer Model Training over Variable-length Sequences

### v1

@@ Appendix B Cost Model
Figure 15: Latency breakdown and fluctuation when training LLaMA 7B on 8 Nvidia A800 GPUs. The left uses pipeline parallel (PP) without data parallel (DP), while the right uses data parallel (DP) without pipeline parallel (PP).
B.1 Memory Model
Consider a model with hidden size $H$, number of layers $\overline{\textup{layers}}$, and vocab size $V$, and there are $N$ available GPUs. During the propagation phase, we utilize the parallel scheme $TP,PP\right>$, while the optimization phase employs a fully sharded strategy. For a sequence length $l$ (i.e., micro-batch shape $[1,l]$), memory usage is calculated as follows. (a) For activations, the memory cost per micro-batch is $\frac{\overline{\textup{layers}}}{PP}\times\frac{l\times H}{TP}\times A$, where $A$ is a constant related to the model architecture, determined through profiling. The cumulative memory peak of all micro-batches (accumulated $PP$ times in the first pipeline) is $\overline{\textup{layers}}\times\frac{l\times H}{TP}\times A$. (b) For model states, given that some are fully sharded while others are related to $TP,PP\right>$, the total memory usage of all layers is $\frac{\overline{\textup{layers}}\times H^{2}}{N}\times B\times\alpha+\frac{\overline{\textup{layers}}}{PP}\times\frac{H^{2}}{TP}\times B\times\left(1-\alpha\right)$. Here, $B$ is a constant related to the model structure, obtainable via profiling, and $\alpha$ is associated with the distribution and precision of model states. For instance, if the optimization phase uses 32-bit optimizer states and parameters, while the propagation phase uses 16-bit parameters and gradients, then $\alpha=\frac{32\times 3}{32\times 3+16\times 2}=\frac{3}{4}$. Similarly, for the embedding table, the memory cost is $\frac{H\times V}{N}\times 16\times\alpha+\frac{H\times V}{TP}\times 16\times\left(1-\alpha\right)$. The total memory peak is the sum of these costs, forming a linear function of $l$. Then, the x-coordinate of this function at the y-coordinate corresponding to GPU memory capacity minus a safety margin provides the maximum sequence length $\texttt{MaxLen}(P)$ supported by $P=\left<TP,PP\right>$. It’s important to note that we can apply a similar modeling approach when incorporating other forms of parallel methods, such as context parallel and sequence parallel.
B.2 Latency Model
@@
Appendix C Enumeration Pruning
@@
Figure 16: (a) LLaMA 13B training efficiency across different parallel schemes, measured in tokens per second. (b) Latency changes pre- and post-sequence packing for two equal-length sequences when $PP=2$, "F" and "B" indicate forward and backward propagation, respectively.
In this section, we will further discuss two trade-offs during packing as a supplement. Based on this, we will perform pruning optimization to avoid some unnecessary enumeration of the number of micro-batches.
  1. 1.
    Less packing, smaller micro-batch size, lower GPU utilization: As shown in Figure 16(a), we assess the GPU’s efficiency in processing sequences of specific lengths by dividing the propagation latency by the number of tokens in the sequence. Our observations indicate that efficiency approaches optimal only when the sequence length exceeds a certain threshold, which is related to intra-op parallel (e.g., TP) but independent of inter-op parallel (e.g., PP). This is due to GEMMs (General Matrix Multiplications) [64]: intra-op parallel reduces the computational load of single matrix multiplication, while inter-op parallel only reduces the number of matrix multiplications. When sequences are short or the intra-op parallel degree is large, the scale of a single matrix multiplication becomes small, leading to under-utilization of the GPU’s SMs and significant kernel launch overhead [63]. Therefore, insufficient packing on some parallel schemes may result in many under-utilized micro-batches, reducing the performance.
  2. 2.
    More packing, smaller number of micro-batches, higher pipeline bubble rate: As shown in Figure 16(b), packing inevitably reduces the number of micro-batches, increasing the bubble rate. In the example shown, there is about a 0.25×0.25$\times$0.25 × performance slowdown. Therefore, packing sequences to the memory limit or the mini-batch maximum is not necessarily optimal.
@@
Appendix D Details of Sequence Dispatching across Pipelines
@@
D.1 Derivation of the Lower Bound
@@ 25             $O_{\text{best}}\leftarrow O_{\text{trial}}$, $M_{\text{best}}\leftarrow\{m_{i,j}\}_{B\times D}$;
D.2 Approximate Solver
[…]
D.3 Performance Comparison
Figure 17: Comparison of estimated propagation latency and actual end-to-end latency between greedy dispatching and ideal dispatching. The experiments use a 32B model on the CommonCrawl dataset, with 64 GPUs. Each iteration handles 200,000 tokens with a maximum sequence length of 32K.
@@ D.3 Performance Comparison
1024 GPUs & 3.2M tokens per iteration 107.21s >1000s 38.56s 13.46s
Table 2: Solving time for each subproblem. The time cost for ideal dispatching, greedy dispatching, and packing is listed for 10 iterations combined. Hydraulis ultimately chooses the greedy dispatching scheme over the ideal dispatching scheme due to its faster speed.
@@
Table 3: Optimal training configurations. DP refers to the data parallel degree, TP to the tensor parallel degree, PP to the pipeline parallel degree, SP to the Ulysses sequence parallel degree, and AC indicates the use of activation checkpointing. For HotSPa, there are 2-3 distinct strategies.
We analyze the gap between our proposed dispatching algorithm, Alg. 1 (referred to as greedy dispatching), and the optimal solution to the original optimization problem, Eq. (3) (referred to as ideal dispatching). As illustrated in Figure 17, we compare the ratios of estimated propagation latency (i.e., the optimal objective given by the algorithm) and actual end-to-end latency by dividing the greedy dispatching results by the ideal dispatching results. The estimations closely align, and the actual end-to-end latency ratios fall between 0.9 and 1.1. This indicates that there is no significant performance difference between the two approaches. Considering the subsequent time analysis in Appendix E, we ultimately opt for greedy dispatching due to its faster solution speed and superior scalability.
Appendix E Time Cost and Scalability
@@
Appendix F Optimal Training Configurations
Table 3 illustrates the manually tuned configurations that achieve the best training performance under different scenarios in § 8.2.

### v2

@@ Appendix C Cost Model
Figure 18. Latency breakdown and fluctuation when training LLaMA 7B on 8 Nvidia A800 GPUs. The left uses pipeline parallel (PP) without data parallel (DP), while the right uses data parallel (DP) without pipeline parallel (PP).
C.1. Memory Model
Consider a model with hidden size $H$, number of layers $\overline{\textup{layers}}$, and vocab size $V$, and there are $N$ available GPUs. During the propagation phase, we utilize the parallel scheme $\left<TP,PP,CP\right$>, while the optimization phase employs a fully sharded strategy. For a sequence length $l$ (i.e., micro-batch shape $[1,l]$), memory usage is calculated as follows. (a) For activations, the memory cost per micro-batch is $\frac{\overline{\textup{layers}}}{PP}\times\frac{l\times H}{TP\times CP}\times A$, where $A$ is a constant related to the model architecture, determined through profiling. The cumulative memory peak of all micro-batches (accumulated $PP$ times in the first pipeline) is $\overline{\textup{layers}}\times\frac{l\times H}{TP\times CP}\times A$. (b) For model states, given that some are fully sharded while others are related to $\left<TP,PP,CP\right$>, the total memory usage of all layers is $\frac{\overline{\textup{layers}}\times H^{2}}{N}\times B\times\alpha+\frac{\overline{\textup{layers}}}{PP}\times\frac{H^{2}}{TP}\times B\times\left(1-\alpha\right)$. Here, $B$ is a constant related to the model structure, obtainable via profiling, and $\alpha$ is associated with the distribution and precision of model states. For instance, if the optimization phase uses 32-bit optimizer states and parameters, while the propagation phase uses 16-bit parameters and gradients, then $\alpha=\frac{32\times 3}{32\times 3+16\times 2}=\frac{3}{4}$. Similarly, for the embedding table, the memory cost is $\frac{H\times V}{N}\times 16\times\alpha+\frac{H\times V}{TP}\times 16\times\left(1-\alpha\right)$. The total memory peak is the sum of these costs, forming a linear function of $l$. Then, the x-coordinate of this function at the y-coordinate corresponding to GPU memory capacity minus a safety margin provides the maximum sequence length $\texttt{MaxLen}(P)$ supported by $P=\left<TP,PP,CP\right$>. It’s important to note that we can apply a similar modeling approach when incorporating other forms of parallel methods, such as context parallel and sequence parallel.
C.2. Latency Model
@@
Appendix D Enumeration Pruning
@@
Figure 19. (a) LLaMA 13B training efficiency across different parallel schemes, measured in tokens per second. (b) Latency changes pre- and post-sequence packing for two equal-length sequences when $PP=2$, “F” and “B” indicate forward and backward propagation, respectively.
In this section, we will further discuss two trade-offs during packing as a supplement. Based on this, we can perform pruning optimization to avoid some unnecessary enumeration of the number of micro-batches.
  1. (1)
    Less packing, smaller micro-batch size, lower GPU utilization: As shown in Figure 19(a), we assess the GPU’s efficiency in processing sequences of specific lengths by dividing the propagation latency by the number of tokens in the sequence. Our observations indicate that efficiency approaches optimal only when the sequence length exceeds a certain threshold, which is related to intra-op parallel (e.g., TP) but independent of inter-op parallel (e.g., PP). This is due to GEMMs (General Matrix Multiplications) (NVIDIA, 2023b): intra-op parallel reduces the computational load of single matrix multiplication, while inter-op parallel only reduces the number of matrix multiplications. When sequences are short or the intra-op parallel degree is large, the scale of a single matrix multiplication becomes small, leading to under-utilization of the GPU’s SMs and significant kernel launch overhead (NVIDIA, 2023a). Therefore, insufficient packing on some parallel schemes may result in many under-utilized micro-batches, reducing the performance.
  2. (2)
    More packing, smaller number of micro-batches, higher pipeline bubble rate: As shown in Figure 19(b), packing inevitably reduces the number of micro-batches, increasing the bubble rate. In the example shown, there is about a $0.25\times$ performance slowdown. Therefore, packing sequences to the memory limit or the mini-batch maximum is not necessarily optimal.
@@
Appendix E Details of Sequence Dispatching across Pipelines
@@
E.1. Derivation of the Lower Bound
@@ 25    $O_{\text{best}}\leftarrow O_{\text{trial}}$, $M_{\text{best}}\leftarrow\{m_{i,j}\}_{B\times D}$;
E.2. Approximate Solver
Our developed approximation algorithm is presented in Alg. 1. We start by randomly shuffling the sequences and then dispatch them one by one (lines 1-6). For each dispatch, we aim to prevent the current slowest pipeline $P_{j}$ from worsening, by considering the accumulated sequence cost $C_{j}$ and extra cost $E_{j}$ (lines 7-14). Specifically, for each potential pipeline $j$ of the current sequence $i$ (line 7), we simulate dispatching the sequence to observe the cost change. We first compute the maximum sequence length based on the current sequence and those already dispatched to the pipeline (line 8). Then, we calculate the new accumulated cost $C_{j}^{\prime}$ by adding the latency of the current sequence $i$ on pipeline $j$ (line 9). After determining the new extra cost $E_{j}^{\prime}$, which is associated with the maximum sequence length and the pipeline parallel (PP) degree (line 10), we evaluate the maximum total cost $O_{\text{max}}$ across all pipelines (line 11). Afterward, we select the pipeline $j^{*}$ that minimizes $O_{\text{max}}$ and update the dispatch matrix and costs accordingly (lines 12-14). This process (i.e., a trial) is repeated $T$ times to select the best result (lines 15-17). [^8] TT is set to 100 in all our experiments. Since multiple trials can be parallelized using multiple processes, the actual execution speed is relatively fast (details in Appendix F), and the gap compared to the optimal solution of Eq. (3) is relatively small (details in Appendix E.3).
E.3. Performance Comparison
Figure 20. Comparison of estimated propagation latency and actual end-to-end latency between greedy dispatching and ideal dispatching. The experiments use a 32B model on the CommonCrawl dataset, with 64 GPUs. Each iteration handles 200,000 tokens with a maximum sequence length of 32K.
Table 2. Solving time for each subproblem. The time cost for ideal dispatching, greedy dispatching, and packing is listed for 10 iterations combined. Hydraulis ultimately chooses the greedy dispatching scheme over the ideal dispatching scheme due to its faster speed.
@@ E.3. Performance Comparison
1024 GPUs & 3.2M tokens per iteration 107.21s ¿1000s 38.56s 13.46s
Table 3. Optimal training configurations. DP refers to the data parallel degree, TP to the tensor parallel degree, $P$ to the pipeline parallel degree, CP to the context parallel degree, SP to the Ulysses sequence parallel degree, and AC indicates the use of activation checkpointing. For HotSPa, there are 2-3 distinct strategies.
@@
Table 4. HotSPa’s switching overhead and iteration time comparison.
Configuration Tokens per Iteration HotSPa Switching Time HotSPa Iteration Time (incl. Switching) Hydraulis Iteration Time
(a) 7B, 32 GPUs 100 1.54s 5.86s 2.21s
(b) 32B, 32 GPUs 100 2.17s 12.21s 7.21s
(c) 32B, 64 GPUs 200 2.21s 13.19s 7.22s
We analyze the gap between our proposed dispatching algorithm, Alg. 1 (referred to as greedy dispatching), and the optimal solution to the original optimization problem, Eq. (3) (referred to as ideal dispatching). As illustrated in Figure 20, we compare the ratios of estimated propagation latency (i.e., the optimal objective given by the algorithm) and actual end-to-end latency by dividing the greedy dispatching results by the ideal dispatching results. The estimations closely align, and the actual end-to-end latency ratios fall between 0.9 and 1.1. This indicates that there is no significant performance difference between the two approaches. Considering the subsequent time analysis in Appendix F, we ultimately opt for greedy dispatching due to its faster solution speed and superior scalability.
Appendix F Time Cost and Scalability
@@
Appendix G Optimal Training Configurations
Table 3 illustrates the manually tuned configurations that achieve the best training performance under different scenarios in §8.2. In terms of HotSPa:
@@ Appendix G Optimal Training Configurations
Appendix H HotSPa’s Switching Overhead
In this section, we analyze the switching overhead of HotSPa under three configurations, as detailed in Table 4. From configuration (a) to (b), we scale the model size, and from (b) to (c), we scale the total training setup, proportionally to LLaMA’s configuration (16M tokens per iteration for 8192 GPUs).
First, when the training workload is relatively small, HotSPa’s switching overhead becomes significant. For example, in configuration (a), 1.54s out of the total 5.86s per iteration (26%) is spent on strategy switching, negatively impacting end-to-end performance.
Second, when scaling from configuration (b) to (c), doubling the data parallel (DP) degrees causes HotSPa’s switching to occur independently within each 32-GPU group. This keeps the switching overhead relatively stable. However, since the tokens per iteration also scale, the overhead remains consistently around 17% of the total iteration time. (Note that our experimental environment is equipped with 400 GB/s NVLink and 200 GB/s IB, which are sufficiently fast.) In contrast, Hydraulis overlaps strategy switching with computation via O-P disaggregation (as evaluated in §8, Table 1), making it a more scalable solution.
Finally, even if the switching overhead were excluded, HotSPa would still lag behind Hydraulis. This is due to the advantages of Hydraulis in terms of strategy optimization and sequence assignment: For one thing, HotSPa only supports homogeneous strategies at any given time, whereas Hydraulis enables dynamic heterogeneous strategies. For another, Hydraulis incorporates fine-grained sequence dispatching and packing, which HotSPa lacks.